# SCBench: A KV Cache-Centric Analysis of Long-Context Methods

### v1

@@ 3.2 Long-Context Shared Context Modes Details
In addition to the carefully designed long-context tasks, we include two shared context modes to more accurately reflect real-world long-context applications: multi-turn mode and multi-request mode, as shown in Fig. 2b.
(i) Multi-turn Mode: A typical scenario in long-context applications, including long-context chat, multi-step reasoning (e.g., Tree-of-Thought (Yao et al., 2024b)), and long-generation CoT. This mode is relevant to long-context methods with KV cache reuse, as the focus in each turn may shift significantly, potentially causing the models to lose information stored in KV cache. Following Zheng et al. (2023a); Wang et al. (2024), we use ground-truth answers instead of model-generated content as the context for follow-up turns.
(ii) Multi-request Mode: Context sharing can occur across sessions or even users, such as multiple users working on the same code repository. In this case, models can encode the shared context and share the memory (KV cache) across multiple requests. Testing long-context methods in such scenarios is crucial, as some require the query for sparse encoding/decoding. For instance, MInference and SnapKV use the final part of the input (often the query) to estimate the overall sparse pattern. This mode tests how well these methods generalize without having the query.
@@ 4 Experiments & Results
We selected six open-source long-context LLMs for our study: Llama-3.1-8B/70B (Dubey et al., 2024), Qwen2.5-72B/32B (Team, 2024), Llama-3-8B-262K (Gradient, 2024), and GLM-4-9B-1M (GLM et al., 2024), along with two gated linear models: Codestal Mamba 7B (team, 2024), and Jamba-1.5-Mini (Lieber et al., 2024). This selection encompasses Transformer, SSMs, and SSM-Attention Hybrid models, representing some of the most effective context lengths among open-source Long-context LLMs. To ensure result stability, all experiments were conducted using greedy decoding in BFloat16 on four NVIDIA A100 GPUs. We evaluated all models using the HuggingFace or vLLM framework with FlashAttention-2 (Dao, 2024) implementation. Additionally, we employed MInference’s implementation (Jiang et al., 2024) to reduce GPU memory overhead. More information of these models and our infrastructure can be found at §C.1.
@@
We evaluated eight categories of long-context solutions on our benchmark, including Gated Linear RNNs (e.g., Codestral-Mamba), SSM-Attention hybrid models (e.g., Jamba), sparse attention, KV cache dropping, prompt compression, KV cache quantization, KV cache retrieval, and KV cache loading, as detailed in Table 1. All methods were tested on Transformer-based long-context LLMs, except for Codestral-Mamba and Jamba. Additionally, we report the KV cache size, pre-filling stage complexity, decoding stage complexity, and whether efficient operations are applied during the pre-filling or decoding stages (details in Sec. 2). The exact implementation and configuration details are provided in §C.2.
@@
Table 4: Average performance of various long-context methods across different base models in two shared context modes on SCBench. For additional results on base models such as Llama-3.1-70B, Qwen2.5-32B, and Llama-3-8B-262K, see Table 10 in §D. Here, $\tau$ denotes the target compression rate.
@@
Table 4, 10, and Fig. 6 illustrate the performance of various long-context methods across multiple tasks and shared context modes in different base LLMs. Key observations include: 1) In retrieval tasks, most long-context methods, except MInference, perform poorly, particularly in exact information retrieval such as string retrieval. 2) Sparse attention methods show significant improvements over sparse decoding methods as the number of request rounds increases, with A-shape demonstrating the greatest enhancement. Tri-shape, which incorporates a dense bottom query tokens into A-shape, boosts first-round performance but has minimal impact on subsequent rounds. Tri-shape also generalizes well across tasks, ranking second only to MInference across models. Our analysis reveals that the Tri-shape bottom improves first-turn instruction-following, thus enhancing overall performance, while A-shape disrupts instruction information, leading to random outputs, as shown in Table 17. 3) KV cache compression methods generally underperform in shared scenarios, showing only slight advantages in the first round. 4) Prompt compression methods enhance global information tasks like many-shot ICL but degrade performance significantly in retrieval-related tasks. 5) SSM-attention hybrid models perform well in single-turn interactions but exhibit degraded accuracy in multi-turn scenarios, especially in RepoQA and Math. Gated Linear RNN models perform poorly in shared context modes.
(a) String Retrieval
(b) Semantic Retrieval
(c) Global Information
@@ 5 Analysis
We analyzed the attention distribution for the Retr.KV task across multiple turns with a shared context. As shown in Fig. 7a, the critical key-value pairs (KVs) are highly query-dependent and vary significantly between turns. We found that, aside from initial and local tokens, attention focuses primarily on the first occurrence of the key in each query. Due to the unpredictability of future queries, the shared context memory (i.e., KV cache in Transformer models) must be fully preserved. This potential distribution shift of importance of history KV states explains why most sub-$O(n)$ decoding methods, particularly KV cache compression methods and pure SSM models, fail in our SCBench benchmark. Previous studies have noted similar issues with SSMs, suggesting that the entire prompt needs to be repeated after each query to recover lost context memory (Arora et al., 2024). In Fig. 7b, we visualize the attention map for the Retr.KV task across turns. While important KVs remain consistent within a turn, they differ significantly between queries. This explains why $O(k)$ KV cache compression methods perform well in single-query tests but fail in follow-up queries. However, the SSM-attention hybrid model Jamba shows potential for reducing overall memory cost by utilizing SSM layers while maintaining $O(n)$ memory in a few attention layers for future lookups (Waleffe et al., 2024). Another promising approach is CPU-GPU collaboration for fast inference, where the full $O(n)$ memory is stored in CPU RAM, and relevant KVs are dynamically loaded to the GPU, achieving sub-$O(n)$ decoding on the GPU (Liu et al., 2024b).
@@
We discussed how sub-$O(n)$ sparse decoding often fails to maintain accuracy across multiple requests with shared context. Interestingly, these sparse approaches perform well in the encoding phase if decoding remains dense. As shown in Fig. 3a, with dense decoding ($O(n)$ memory), Tri-Shape and A-Shape demonstrate strong performance in multi-request testing. While this success of sparse encoding with dense decoding has been observed in single-turn tests (Sun et al., 2024c; Jiang et al., 2024), we are the first to showcase its potential in shared context scenarios. In contrast, extending sparse patterns to the decoding stage leads to significant performance degradation (e.g., StreamingLLM). Even with dense encoding, sparse decoding methods generally underperform in shared context testing, particularly KV cache compression methods. This disparity may be due to redundancy in the encoding output, while decoding plays a critical role in generation tasks (Deng et al., 2024). Redundant input prompts allow key information to be captured even with sparse encoding, but sparse decoding reduces per-layer connectivity, limiting the model’s ability to focus on critical tokens. Since sparse decoding relies on proxy tokens for global information access, it restricts the construction of complex attention functions (Yun et al., 2020). We emphasize the need for more sophisticated sparse patterns in sparse attention. Dynamic sparse attention methods can enhance connectivity and enable faster information propagation (Jiang et al., 2024), better approximating full attention performance compared to static sparse patterns, as shown in Fig. 6.
@@
While $O(n)$ memory is essential in multi-request scenarios with shared context, this requirement can be relaxed for highly compressible inputs in simpler tasks. For instance, the Needle-in-the-Haystack benchmark (Kamradt, 2023) embeds key information (the "needle") into repetitive noise (the "haystack"), allowing sub-$O(n)$ methods to achieve reasonable accuracy since the noise is highly compressible. Similarly, tasks like summarization involve compressible contexts, where sub-$O(n)$ methods can balance efficiency and performance. However, with dynamic and complex inputs, sub-$O(n)$ methods often fail to store all necessary information, resulting in poor performance on challenging retrieval tasks. Tasks like Retr.KV and Retr.Prefix-Suffix, which involve random and incompressible key-value pairs and strings, require models to fully utilize their context window. In summary, while compressible tasks may overestimate a model’s capabilities, sub-$O(n)$ methods remain useful for simpler tasks due to their efficiency.
@@
One concern with long-context methods in KV cache reuse scenarios is their reliance on the query for compression to achieve efficient encoding or decoding. However, in real-world applications, a single context is often shared across multiple queries, requiring these methods to function without the query. This raises the question: can query-dependent long-context methods generalize effectively without it? In Table 5, we compare the performance of three query-awareness long-context methods w/ and w/o the query provided, highlighting degraded performance in the absence of the query using underlines. We observed that both the KV cache compression method SnapKV and the static sparse attention method Tri-shape struggled to maintain accuracy without the query. In contrast, the dynamic sparse attention method MInference demonstrated more robust generalization, likely due to its dynamic and sophisticated sparse patterns, particularly the presence of diagonal connections in its attention map.
6 Related Works
@@
Long-context inference faces two key bottlenecks: computational cost during pre-filling and memory cost during decoding (Fu, 2024). Pre-filling optimizations include state space models (Gu & Dao, 2024; Gu et al., 2022), linear attention methods (Peng et al., 2023; Sun et al., 2023), memory-based approaches (Munkhdalai et al., 2024), sparse attention (Jiang et al., 2024), hybrid techniques (Lieber et al., 2024; Ho et al., 2024; Ren et al., 2024; Xiao et al., 2024a), and prompt compression (Li et al., 2023b; Jiang et al., 2023; Pan et al., 2024). Decoding optimizations focus on: 1) Attention KV reuse to reduce storage (Shazeer, 2019; Ainslie et al., 2023; Sun et al., 2024c; Liu et al., 2024a; Nawrot et al., 2024); 2) Static KV compression (Xiao et al., 2024b; Han et al., 2024); 3) Dynamic KV compression, including cache discarding (Zhang et al., 2024b; Ge et al., 2024; Liu et al., 2023; Li et al., 2024c) and offloading (Ribar et al., 2024; Tang et al., 2024; Dai et al., 2024; Liu et al., 2024b; Chen et al., 2024; Sun et al., 2024a); 4) Hierarchical speculative decoding (Sun et al., 2024b). Most methods are tested on single-turn benchmarks and employ query-conditioned lossy techniques, which may degrade performance in multi-turn scenarios with prefix caching. This limitation motivates the design of SCBench, a benchmark that evaluates long-context solutions in shared context settings.
7 Conclusion
This paper addresses a key gap in the evaluation of long-context methods, which used to fully rely on single-turn interactions and overlook the performance of shared long-context scenarios that is common in real-world LLM applications. To bridge this gap, we introduce SCBench, a comprehensive benchmark designed to test long-context methods with KV cache reuse across multiple domains, featuring 12 tasks that span four long-context capabilities: string retrieval, semantic retrieval, global information processing, and multi-tasking, across two shared context modes. Using our benchmark, we category long-context methods into 4 stage in a KV-cache centric perpective, including KV cache generation, compression, retrieval, loading. We evaluated eight categories of long-context methods, including gated linear RNNs, hybrid models, sparse attention, KV cache dropping, quantation, retrieval, loading, and prompt compression, on eight state-of-the-art LLMs, including Llama-3.1-8B/70B, Qwen2.5-72B/32B, Llama-3-8B-262K, GLM-4-9B, Codestal Mamba, and Jamba-1.5. Our results show a clear disparity in KV cache management: methods maintaining KV cache at $O(n)$ excel in multi-request scenarios, while sub-$O(n)$ methods perform well in single-turn settings but struggle with complex interactions. These findings highlight the importance of multi-turn, shared-context scenarios in developing and evaluating long-context methods, offering a more realistic benchmark and key insights for improving long-context models and future architecture design in practical applications.
Appendix A Compared to Prior Long-Context Benchmark
@@
Appendix B Hyper-Parameters of Efficient Long-Context Methods
@@
Appendix C Experiment Details
C.1 Long-context Methods Details
@@
interleaves blocks of Transformers and Mamba layers, aiming to obtain the benefits of both architecture, i.e., the expressive power of Transformer and the linear complexity of Mamba layers. Jamba (Lieber et al., 2024) and Samba (Ren et al., 2024) are representative efforts on this direction. Waleffe et al. (2024) also highlights the potential of such hybrid architectures and found only a few number of attention layers can lead to significant performance increase compared to pure SSMs models.
@@
is extensively studied for long sequence processing, including image synthesis and multi documents question answering. We test three sparse attention approach in our paper: A-shape, Tri-shape, and MInference. In A-shape attention, each token is only allowed in to attend to initial tokens and local tokens, resulting a A-shape on its attention map (Xiao et al., 2024b). Tri-shape attention is a variant of A-shape method, we introduced in our paper, where we add a dense attention space at the bottom of the triangle A-shape attention matrix. This is based on the promising results of sparse encoding with dense decoding, where the dense space we added is a natural extrapolate of the dense decoding idea. MInference (Jiang et al., 2024) is the state-of-the-art dynamic sparse attention approach where the exact sparse pattern are dynamically built on-the-fly to better approximate full attention operation.
@@
is a series of studies that attempt to solve the linearly growing memory (often referred as KV Cache) cost in LLMs inference. For example, StreamingLLM (Xiao et al., 2024b) use a constant size of KV Cache in their decoding steps, where only the state of initial and local tokens are preserved, and the rest part of KV Caches are evicted from the memory. SnapKV (Li et al., 2024c) introduces the concept of the observation window. It selects the top-K KVs that are extensively attended to in the observation window, and removes other KVs from the Cache. This method was reported to performance well in simple Neeld-in-A-Haysatck tasks and many other natural language tasks.
@@
C.2 Additional Implementation Details

### v2

@@ 3.2 Long-Context Shared Context Modes Details
In addition to the carefully designed long-context tasks, we include two shared context modes—multi-turn and multi-request—to better reflect real-world applications, as shown in Fig. 2b.
(i) Multi-turn Mode: A typical scenario in long-context applications includes long-context chat, multi-step reasoning (e.g., Tree-of-Thought (Yao et al., 2024b)), and long-generation CoT. This mode is relevant for long-context methods with KV cache reuse, as focus shifts across turns can lead to information loss in the KV cache. Following Zheng et al. (2023a); Wang et al. (2024), we use ground-truth answers instead of model-generated content as the context for follow-up turns.
(ii) Multi-request Mode: Context sharing spans sessions or users, such as collaborators on a shared code repository. Models can encode shared context and reuse the KV cache across requests. Evaluating long-context methods is crucial, as some depend on the query for sparse encoding/decoding. For instance, MInference and SnapKV use the input’s final segment (typically the query) to estimate the sparse pattern, assessing their generalization without query access.
@@ 4 Experiments & Results
We selected six open-source long-context LLMs: Llama-3.1-8B/70B (Dubey et al., 2024), Qwen2.5-72B/32B (Team, 2024), Llama-3-8B-262K (Gradient, 2024), and GLM-4-9B-1M (GLM et al., 2024), along with two gated linear models: Codestal Mamba 7B (team, 2024) and Jamba-1.5-Mini (Lieber et al., 2024). This selection covers Transformer, SSMs, and SSM-Attention Hybrid models, representing leading open-source long-context LLMs. For stability, all experiments used greedy decoding in BFloat16 on four NVIDIA A100 GPUs. We evaluated models via HuggingFace or vLLM with FlashAttention-2 (Dao, 2024) and leveraged MInference (Jiang et al., 2024) to reduce GPU memory overhead. More details on these models and infrastructure are in §D.1.
@@
We evaluated eight long-context solution categories on our benchmark: Gated Linear RNNs (e.g., Codestral-Mamba), SSM-Attention hybrids (e.g., Jamba), sparse attention, KV cache dropping, prompt compression, KV cache quantization, retrieval, and loading, as detailed in Table 1. All methods were tested on Transformer-based long-context LLMs, except Codestral-Mamba and Jamba. We also report KV cache size, pre-filling and decoding complexity, and whether efficient operations are applied (details in §2). More details are shown in §D.2.
@@
Table 4: Average performance of various long-context methods across different LLMs in two shared context modes on SCBench. For additional results on base models such as Llama-3.1-70B, Qwen2.5-32B, and Llama-3-8B-262K, see Table 10 in §E. Here, $\tau$ denotes the target compression rate.
@@
Tables 4, 10, and Fig. 9 present the performance of various long-context methods across tasks and shared context modes in different base LLMs. Key findings include: 1) In retrieval tasks, most methods, except MInference, perform poorly, especially in exact retrieval tasks like string matching. 2) Sparse attention outperforms sparse decoding as request rounds increase, with A-shape showing the greatest improvement. Tri-shape, which adds dense bottom query tokens to A-shape, enhances first-round performance but has little effect on later rounds. It generalizes well across tasks, ranking second only to MInference. Our analysis indicates that Tri-shape improves first-turn instruction-following, boosting overall performance, while A-shape disrupts instruction information, causing random outputs (Table 17). 3) KV cache compression methods generally underperform in shared scenarios, offering only minor benefits in the first round. 4) Prompt compression improves global information tasks like many-shot ICL but significantly degrades retrieval-related performance. 5) SSM-Attention hybrids perform well in single-turn interactions but suffer accuracy drops in multi-turn tasks, especially in RepoQA and Math. Gated Linear RNN models struggle in shared context modes.
@@ 5 Analysis
Fig. 6b visualizes the attention map for the Retr.KV task across turns. While important KVs remain stable within a turn, they vary significantly between queries. This explains why $O(k)$ KV cache compression methods perform well in single-query tests but fail in follow-ups. However, the SSM-attention hybrid model Jamba shows promise in reducing memory costs by using SSM layers while maintaining $O(n)$ memory in select attention layers for future lookups (Waleffe et al., 2024). Another promising approach is CPU-GPU collaboration, where full $O(n)$ memory is stored in CPU RAM, dynamically loading relevant KVs to the GPU for sub-$O(n)$ decoding (Liu et al., 2024b).
@@
We examined how sub-$O(n)$ sparse decoding struggles to maintain accuracy across multiple requests in shared context scenarios. Interestingly, sparse methods perform well in encoding if decoding remains dense. As shown in Fig. 3a, with dense decoding ($O(n)$ memory), Tri-Shape and A-Shape achieve strong multi-request performance. While this phenomenon has been noted in single-turn tests (Sun et al., 2024c; Jiang et al., 2024), we are the first to demonstrate its potential in shared context settings. Conversely, extending sparse patterns to decoding severely degrades performance (e.g., StreamingLLM). Even with dense encoding, sparse decoding methods, particularly KV cache compression, perform poorly in shared context scenarios. This may stem from redundancy in encoding outputs, whereas decoding plays a crucial role in generation (Deng et al., 2024). Sparse encoding can still capture key information due to redundant input prompts, but sparse decoding weakens per-layer connectivity, limiting focus on critical tokens. Since sparse decoding relies on proxy tokens for global access, it constrains the formation of complex attention functions (Yun et al., 2020). We highlight the need for more advanced sparse patterns in sparse attention. Dynamic sparse attention can improve connectivity and accelerate information propagation (Jiang et al., 2024), better approximating full attention performance compared to static sparse patterns (Fig. 9).
@@
While $O(n)$ memory is crucial for multi-request scenarios with shared context, it can be relaxed for highly compressible inputs in simpler tasks. For example, the Needle-in-the-Haystack benchmark (Kamradt, 2023) embeds key information (the "needle") within repetitive noise (the "haystack"), allowing sub-$O(n)$ methods to achieve reasonable accuracy due to the high compressibility of noise. Similarly, summarization tasks involve compressible contexts, enabling sub-$O(n)$ methods to balance efficiency and performance. However, for dynamic and complex inputs, sub-$O(n)$ methods often fail to retain all necessary information, leading to poor performance in challenging retrieval tasks. Tasks like Retr.KV and Retr.Prefix-Suffix, which involve random and incompressible key-value pairs and strings, require models to fully utilize their context window. In summary, while compressible tasks may overestimate a model’s capabilities, sub-$O(n)$ methods remain viable for simpler tasks due to their efficiency.
@@
A key concern with long-context methods in KV cache reuse scenarios is their reliance on the query for compression to enable efficient encoding or decoding. However, in real-world applications, a single context is often shared across multiple queries, requiring these methods to operate without query access. This raises the question: Can query-dependent long-context methods generalize effectively without it? Table 5 compares the performance of three query-aware long-context methods with and without the query, highlighting performance degradation in its absence (underlined). We found that both the KV cache compression method SnapKV and the static sparse attention method Tri-Shape struggled to maintain accuracy without the query. In contrast, the dynamic sparse attention method MInference exhibited stronger generalization, likely due to its adaptive sparse patterns, particularly its diagonal connections in the attention map.
6 Conclusion
This paper addresses a key gap in evaluating long-context methods, which have traditionally focused on single-turn interactions while overlooking shared long-context scenarios—common in real-world LLM applications. To bridge this, we introduce SCBench, a comprehensive benchmark assessing long-context methods with KV cache reuse across 12 tasks, covering string retrieval, semantic retrieval, global information processing, and multi-tasking, evaluated in two shared-context modes. Using our benchmark, we categorize long-context methods into four KV cache-centric stages: generation, compression, retrieval, and loading. We evaluate eight method categories (e.g., gated linear RNNs, hybrid models, sparse attention, KV cache dropping, quantization, retrieval, loading, and prompt compression) on eight state-of-the-art LLMs, including Llama-3.1-8B/70B, Qwen2.5-72B/32B, Llama-3-8B-262K, GLM-4-9B, Codestal Mamba, and Jamba-1.5. Our results reveal a clear KV cache management trade-off: $O(n)$ methods excel in multi-request scenarios, while sub-$O(n)$ methods perform well in single-turn settings but struggle with complex interactions. These findings highlight the need for evaluating long-context methods in shared-context, multi-turn scenarios, offering a more realistic benchmark and valuable insights for improving future long-context models and architectures.
Appendix A Related Works
@@
Long-context inference faces two key bottlenecks: computational cost during pre-filling and memory cost during decoding (Fu, 2024). Pre-filling optimizations include state space models (Gu & Dao, 2024; Gu et al., 2022), linear attention methods (Peng et al., 2023; Sun et al., 2023), memory-based approaches (Munkhdalai et al., 2024), sparse attention (Jiang et al., 2024; Zhu et al., 2024; Lai et al., 2025; Yuan et al., 2025; Lu et al., 2025), hybrid techniques (Lieber et al., 2024; Ho et al., 2024; Ren et al., 2025; Fu et al., 2024; Xiao et al., 2025; Yang et al., 2025), and prompt compression (Li et al., 2023b; Jiang et al., 2023; Pan et al., 2024). Decoding optimizations focus on: 1) Attention KV reuse to reduce storage (Shazeer, 2019; Ainslie et al., 2023; Sun et al., 2024c; Liu et al., 2024a; Nawrot et al., 2024); 2) Static KV compression (Xiao et al., 2024; Han et al., 2024); 3) Dynamic KV compression, including cache discarding (Zhang et al., 2024b; Ge et al., 2024; Liu et al., 2023; Li et al., 2024c) and offloading (Ribar et al., 2024; Tang et al., 2024; Dai et al., 2024; Liu et al., 2024b; Chen et al., 2025; Sun et al., 2024a; Hooper et al., 2024; Desai et al., 2024); 4) Hierarchical speculative decoding (Sun et al., 2024b). Most methods are tested on single-turn benchmarks and employ query-conditioned lossy techniques, which may degrade performance in multi-turn scenarios with prefix caching. This limitation motivates the design of SCBench, a benchmark that evaluates long-context solutions in shared context settings.
Appendix B Compared to Prior Long-Context Benchmark
@@
Appendix C Hyper-Parameters of Efficient Long-Context Methods
@@
Appendix D Experiment Details
D.1 Long-context Methods Details
@@
interleaves blocks of Transformers and Mamba layers, aiming to obtain the benefits of both architecture, i.e., the expressive power of Transformer and the linear complexity of Mamba layers. Jamba (Lieber et al., 2024) and Samba (Ren et al., 2025) are representative efforts on this direction. Waleffe et al. (2024) also highlights the potential of such hybrid architectures and found only a few number of attention layers can lead to significant performance increase compared to pure SSMs models.
@@
is extensively studied for long sequence processing, including image synthesis and multi documents question answering. We test three sparse attention approach in our paper: A-shape, Tri-shape, and MInference. In A-shape attention, each token is only allowed in to attend to initial tokens and local tokens, resulting a A-shape on its attention map (Xiao et al., 2024). Tri-shape attention is a variant of A-shape method, we introduced in our paper, where we add a dense attention space at the bottom of the triangle A-shape attention matrix. This is based on the promising results of sparse encoding with dense decoding, where the dense space we added is a natural extrapolate of the dense decoding idea. MInference (Jiang et al., 2024) is the state-of-the-art dynamic sparse attention approach where the exact sparse pattern are dynamically built on-the-fly to better approximate full attention operation.
@@
is a series of studies that attempt to solve the linearly growing memory (often referred as KV Cache) cost in LLMs inference. For example, StreamingLLM (Xiao et al., 2024) use a constant size of KV Cache in their decoding steps, where only the state of initial and local tokens are preserved, and the rest part of KV Caches are evicted from the memory. SnapKV (Li et al., 2024c) introduces the concept of the observation window. It selects the top-K KVs that are extensively attended to in the observation window, and removes other KVs from the Cache. This method was reported to performance well in simple Neeld-in-A-Haysatck tasks and many other natural language tasks.
@@
D.2 Additional Implementation Details